Project

General

Profile

NDOS Commissioning Meeting 2011/05/02

Attendance: Jon, Enrique, Mat, Kurt, Tom, Athans, Denis, Mark Mathis, Nate, Sue, Jarek, Jan, Nick,

Agenda:

  1. NDOS Status/Installation Plans
  2. Readiness for HE Beam
  3. NDOS Photos
  4. Proof of Principle of FEB Readout in ACNET
  5. DCS Plans
  6. DataCheck update
  7. DAQ/Electronics Update & Plans
  8. AOB

NDOS Status/Installation Plans

  • Mat: there was a hold up on the scint. top-off last week because of tanker; but the qualified transfer line got contaminated (mud in hose). This is being worked on, hopefully top-off will resume sometime later this week.
  • Denis (cap-board installation): first block is ~complete (2 FEBs remain). 2nd block: installation on top just began, will resume tomorrow. Estimate to finish 1st di-block is less than a day. A couple of more days to complete everything. Denis is working on quantifying the install. success rate. So far, only 40% are "good". In general we are seeing an impact, but noise is still removed. Denis will post some plots.
  • Mat: want to put 10 APDs on top of detector; will schedule the work for later this week, or possibly today while the MAP is free

Readiness for HE Beam

  • NuMI is now running in LE beam configuration (as of 10 am today). NuMI will run in HE beam config. once NT-01 (NuMI Target 01) is ready to be reinstalled; current time estimate for this is ~2 weeks. The duration of the HE run is expected to be ~1 week.
  • Tricia has requested that we collect as much HE beam data as possible so we can use these data for normalization. Jon has proposed that all disruptive installation work be put on hold during the HE run in order to maximize the amount of useful data we collect.
  • Mark Messier is unconvinced we should postpone the installation work. What do we really gain from the HE data? Jon and Mark will ask Tricia to elaborate.

NDOS Photos

  • Professional photos of the NDOS by FNAL Media Services is scheduled for 2 pm tomorrow (Tuesday, May 3, 2011).
  • Please help cleanup the area if you have any junk lying around.

Proof of Principle of FEB Readout in ACNET

  • Document written by Dennis Nicklaus, AD Controls Dept.
  • Jon: how will a decision be made on the two approaches (EPICS vs. ACNET). Athans: will depend on scalability, readiness
  • Jon: next step has to be trying to run these systems on the test_stand
  • Andrew: we've promised the FNAL people working on this a requirements document. We need to do this ASAP. Athans: just began working on this. Andrew: others can help.
  • Kurt: will produce some notes on design issues, how to integrate with DAQ, etc.
  • Denis: pedestalRunner needs to be part of the design

DCS Plans

  • Athans will communicate with DAQ group and will work on requirements document
  • Jon will follow up with Chad later

DataCheck update

  • Jarek presented
  • Mark: it would be good to have DataCheck produce a list of good runs and subruns

DAQ/Electronics Update & Plans

  • Andrew: TDU was moved to MINOS. Accel. signals were hooked in last Tuesday. We haven't had any problems syncing, but we are having problems booting the PPC remotely and running the SpillServer. The goal is get a SpillServer running on that TDU today. This will be a redundant system. But the GlobalTrigger software needs to be able to decide which server to listen to.
  • Mat: DCM 4-2? Sue has looked at the log output from Mat and Denis's tests. It is odd that the tests that failed don't involved the FPGAs. Sue believes there is something wrong with the hardware. Possibly the network connection. Another possibility is CPU or memory. Suggests we swap it out with one from the test-stand. dcm-01 might be available. Mat will pick up whichever DCM from the test-stand and swap it with 4-2.

AOB